A line L1 passes through point[1,2]and has a gradient of 5.Another line L2 is perpendicular to L1 and meets it at Point x=4 .find the equation of L2 in the form y=mx+c
Answer by mananth(16946)

y-2=5(x-1)
y-2 =5x-5
y=5x-3
Slope of perpendicular line will be (-1/5)
Passes through (4,0)
(y-y1)= m(x-x1)
(y-0)=(-1/5)(x-4)
y= (-1/5)x+(4/5)